# LIST: Learning to Index Spatio-Textual Data for Embedding based Spatial Keyword Queries

### 12 Mar 2024 v1

@@ III-A Preliminary
We consider a geo-textual object dataset $D$, where each geo-textual object $o\in D$ has a location description $o.loc$ and a textual document description $o.doc$. The location description $o.loc$ is a two-dimensional GPS coordinate composed of latitude and longitude. The document description $o.doc$ reflects the textual content. The Top-k KNN Spatial-Keyword Query (TkQ[2] is defined as follows.
@@
III-B Data Analysis and Motivations
@@
Word Mismatch. As discussed earlier, existing indexes [2, 7, 9] utilize traditional ranking methods, such as BM25 [10] and TF-IDF [11], to compute textual relevance. These models often suffer from the word mismatch issue, which reduces their effectiveness. As shown in Figure 1(a), in the Beijing dataset, nearly 20% of the ground-truth query-object positive (relevant) pairs have no overlap of words. Since these traditional models rely on exact word matching to compute text relevance, relevant objects may receive a low or even zero textual relevance from these models. For instance, given a search ‘nearby drugstore’, if the search engine employs the BM25 to determine textual relevance, a drugstore labeled as ‘pharmacy’ would get a textual relevance score of zero.
@@
Spatial Relevance. Previous studies [2, 5] typically compute spatial relevance by $1-SDist(q.loc,o.loc)$. The implicit assumption for the linear formulation is that the user preference for geo-textual objects is a linear function of distance. However, this assumption was not examined previously and we found it does not really hold on our real-life datasets. In Figure 1(b), we illustrate this issue by comparing the cumulative distribution function (CDF) of ground-truth positive (relevant) query-object pairs with the ‘Linear Distribution’, a linearly ascending hypothetical scenario that positive query-object pairs are uniformly distributed across the range $[0,1]$. This figure shows a sharp increase in the CDF of ground-truth positive pairs for $SDist(q.loc,o.loc)$ below 0.1, a pattern that greatly differs from the hypothetical scenario.
@@
Efficiency. To address this issue, several recent studies [5, 13, 57] have utilized deep learning techniques to evaluate textual relevance, thereby enhancing ranking effectiveness. However, these methods rely on word embeddings and deep neural networks to compute textual relevance, resulting in high querying latency. For instance, DrW [5] needs to identify the top-k relevant words from each object for each query word, and then aggregate these word pairs using an attention mechanism to compute the textual relevance. On the Geo-Glue dataset, which comprises 2.8 million objects, DrW takes more than 7 seconds to answer a query on average in our experiment, which aligns with the results reported in [5]. This makes them unsuitable as a retriever for practical geo-textual object retrieval applications, although they can be used as rerankers for a small number of retrieved objects.
@@
Incorporating deep textual relevance into spatial keyword queries has presented significant efficiency challenges. On the one hand, the index techniques developed for spatial keyword queries are designed based on the traditional retrieval models such as BM25 and cannot be used to handle embedding-based spatial keyword queries. On the other hand, existing ANNS indexes, which are discussed in Section II, do not consider spatial relevance. It is an open problem of incorporating spatial relevance into the ANNS indexes to support spatial keyword queries. Directly employing existing ANNS indexes for spatial keyword queries will result in severe degradation of effectiveness (shown in Section V-B).
III-C Problem Statement
To address these issues, we aim to (1) design a lightweight and effective embedding-based spatio-textual relevance model as the ranking function, and (2) develop an index that can accelerate the top-k search process of embedding based spatial keyword queries.
IV PROPOSED RETRIEVER (LIST)
@@
IV-A Overview
As discussed in the introduction, LIST is featured with a novel machine learning based ANNS index and a new relevance model. Notably, our index is applicable to any textual relevance model that transforms query keywords and object text descriptions into two separate embeddings.
The workflow of LIST is illustrated in Figure 2, which has three phases: the training, indexing, and query phase. During the training phase, we first train our relevance model, and then train our index. During the indexing phase, each object is partitioned into one cluster. During the query phase, the trained index routes each query to a cluster or a subset of clusters that have the highest probabilities. In the cluster, LIST returns $k$ objects with the highest scores ranked by the trained relevance model as the query result.
IV-B The Proposed Relevance Model
Textual Relevance Learning. Inspired by the dual-encoder model’s great success in document retrieval [38], we employ a dual-encoder module to encode both the query keywords and the document into two separate embeddings and calculate textual relevance using the two embeddings, which is used in spatial keyword query for the first time. Compared with previous relevance models [5] which rely on word embeddings and complex nerual networks, this approach is more efficient.
This module comprises an object encoder $E_{o}$ and a query encoder $E_{q}$, which are pre-trained language models like Bert [58]. The encoder takes the textual content of a query or object as input, captures interactions between terms by a transformer-based model, and finally utilizes the representation of the [CLS] token as the global semantic representation, which is a $d$-dimensional dense vector. This process is formulated as:
@@ IV-B The Proposed Relevance Model
where $E_{o}(\textperiodcentered;\theta_{o})$ denotes the object encoder parameterized with $\theta_{o}$ and $E_{q}(\textperiodcentered;\theta_{q})$ denotes the query encoder parameterized with $\theta_{q}$. Then the text relevance score can be calculated by the inner product between the $q.emb$ and $o.emb$,
@@
Spatial Relevance Learning. To fit the real-world scenario described in Section III-B, it is essential to assign higher spatial relevance to nearby objects and lower spatial relevance to distant objects. Therefore, we propose a novel learning-based spatial relevance module to better capture spatial relevance, where $S_{in}=1-SDist(q.loc,o.loc)$ is taken as input.
The user’s preference for geo-textual objects increases as the distance decreases and exhibits a stepwise decline as the distance increases, i.e., users tend to visit nearby objects and become less tolerant as the distance increases. To illustrate this stepwise pattern, let us consider a scenario in which a customer wishes to purchase coffee from Starbucks. If the nearest Starbucks is very close, s/he would go and buy coffee. If the nearest Starbucks is a little far, s/he may hesitate and her/his intention of visiting the Starbucks would decrease. If the nearest Starbucks is very far away, s/he would give up this idea. Therefore, a new spatial relevance module is designed, which is a learnable monotonically increasing step function, to align with this real-world characteristic.
Our learning-based spatial relevance module consists of a threshold array $\mathcal{T}\in\mathbb{R}^{t\times 1}$ and a learnable weight array $w_{s}\in\mathbb{R}^{1\times t}$. Here, $\mathcal{T}$ is used to ensure that the learned function is a step function and $w_{s}$ is employed to learn and estimate the spatial relevance. Specifcially, $\mathcal{T}$ stores the threshold values that determine the transition points of the step function, i.e., the value exceeded by $S_{in}$ will trigger an increase of spatial relevance. $\mathcal{T}$ is structured as $\mathcal{T}[i]=\frac{i}{t}$, where $i\in[0,t]$ and $t$ is a hyperparameter to control the increment of the threshold value. For example, when $t=100$, $\mathcal{T}$ is $[0.0,0.01,\cdots,0.99,1.0]$. The learnable weight array $w_{s}$ determines the extent of the increase when the input $S_{in}$ reaches these threshold values. Note that weights in $w_{s}$ are learned from the training data. When the input $S_{in}$ exceeds the value of $\mathcal{T}[i]$, then the spatial relevance increases by $act(w_{s}[i])$, where $act$ is an activation function to ensure $act(w_{s}[i])$ remain non-negative. This process ensures that the output $SRel$ exhibits a step increase as the input $S_{in}$ increases. The learned spatial relevance is computed as below:
@@
Adaptive Weight Learning. The latest study [5] has showed the importance of weight learning in improving ranking effectiveness. We rely on the dual-encoder architecture to obtain embedding of query $q.emb$ to compute textual relevance. In addition, our empirical experiments show that the location embedding technique would introduce additional computational costs. Hence, we employ a simple yet effective manner that directly utilizes an MLP layer to determine an optimal weight based on $q.emb$, formulated as follows:
@@
Training Strategy. To train our relevance model, we employ the contrastive learning strategy. Given a query $q_{i}$, the positive (relevant) geo-textual object $o_{i}^{+}$ is obtained by real-world ground-truth data (detailed in Section V-A). For the negative (irrelevant) geo-textual objects, we choose a subset of hard negative objects for training. Specifically, following previous work [5], for each training query, in each training epoch, we randomly pick $b$ hard negative samples from this set. We optimize the loss function as the negative log-likelihood of the positive object:
@@
In addition, we also utilize the in-batch negatives strategy [59] to further enhance training efficiency.
IV-C The Proposed Index
Figure 3 illustrates our proposed index, which takes geo-location and textual embeddings as input. It uses a Multi-Layer Perception (MLP) to produce the $c$-cluster probability distribution and employs training strategies to ensure relevant queries and objects have similar probability distributions.
@@ IV-C The Proposed Index
Here $x$ is the input feature vector for the cluster classifier. Although several region embedding techniques exist [60], we find that this simple approach is sufficiently effective. Cluster Classifier. Our cluster classifier is a lightweight Multi-Perception Layer (MLP). Given $x$ as the input, it produces the $c$-cluster probability distribution, defined as follows:
@@
where $c$ is a hyperparameter that indicates the desired number of clusters we aim to obtain, and $Prob$ represents the predicted $c$-cluster probability distribution. The MLP is shared between queries and objects. The $c$-cluster probability distribution of object $o$ is denoted as $Prob_{o}$ and that of query $q$ is represented as $Prob_{q}$. The cluster classifier actually serves as a relevance classifier, grouping identified relevant queries and objects into the same cluster. For example, as illustrated in Figure 3, the object $o_{4}$ and query $q_{2}$ forms a positive query-object pair, and they are expected to be grouped into the same cluster $C_{2}$.
Figure 3: The illustration of the index.
Pseudo-Label Generation. To train the cluster classifier, both positive and negative pairwise labels are required. We utilize the ground-truth query-object relevant label as the positive pairwise label for training. However, we lack high-quality negative pairwise labels. Randomly selecting negative objects will lead to overfitting and all objects are grouped together (as detailed later). Therefore, we propose a novel pseudo pairwise negative label generation method.
@@
This adaptive pseudo-negative generation method draws inspiration from hard negative sample strategy [44, 61]. What sets our approach apart from existing studies is that we employ an adaptive manner to select hard negative samples, thereby controlling the difficulty level of the generated negative samples. This adjustment strikes the trade-off between the effectiveness and efficiency.
@@
On the other hand, when our index is trained on a set of harder negative objects, it might also exclude some positive ones and reduce its effectiveness. Thus, the choice of $neg_{start}$ strikes a trade-off between effectiveness and efficiency. This alleviates the lack of negative label issue and the skewed cluster distribution issue of existing techniques [18, 19].
Training Strategy. Based on the hard negative objects and positive objects provided above, we employ the MCL loss function [62] to train the MLP, as described below:
@@
As we aim at grouping the relevant pairs of query and object into the same cluster while the irrelevant pairs are into distinct clusters, the queries and objects are jointly incorporated. This is because the MLP is shared between objects and queries. In this manner, the complex relationships between queries and objects can be well learned in the process of index construction. Figure 3 also demonstrates this process, where relevant pairs share similar output and irrelevant ones have different output.
Learning to Partition and Route. During the indexing phase, each object $o$ is partitioned to the cluster with the highest probability according to $Prob_{o}$. Once the partitioning is completed, the objects assigned to a cluster are stored in a corresponding list, which can act as an inverted file for these objects. Each object in this list is represented by a $d$-dimensional vector $o.emb$, and a geo-location $o.loc$, which are utilized to calculate the relevance score for incoming queries. During the query phase, a given query $q$ is directed to the cluster that has the highest probability. Subsequently, the relevance scores between $q$ and all objects in the cluster are calculated and the top-k objects are selected as result objects.
@@
Insertion and Deletion Policy. When a new object comes, we convert it into an embedding using the trained relevance model and then assign it to specific clusters using the trained index. When an object is deleted, we simply remove it from the corresponding cluster. Note that most existing ANNS indexes are static [48, 16, 17]. When inserted data significantly differs in distribution from the existing data or when insertions occur frequently, periodically rebuilding the index is necessary to maintain high accuracy. Our index follows this approach. Apparently, the time costs of insertion and deletion operations are negligible.
Cluster Evaluation. After the training phase, we employ the trained index to produce clusters $C=\{C_{1},C_{2},\cdots,C_{c}\}$, where $C_{i}$ represents cluster $i$. We use validation queries to evaluate the quality of the clusters. The validation queries are fed into the trained index and routed to a cluster. $C_{i}^{q}$ denotes a list containing the validation queries routed to $C_{i}$, $C_{i}^{o}$ denotes a list containing objects partitioned to $C_{i}$, and ||||| | indicates the size of a list. We proceed to introduce two metrics to evaluate the quality of the clusters.
@@
In addition, we also consider concerns of efficiency. For this purpose, we introduce another metric, the Imbalance Factor (IF) [63], which measures the degree of balance across all clusters, denoted as $\text{IF}(C)$. The $IF(C)$ is formulated as:
@@
where $\text{IF}(C)$ is minimized when $|C_{1}|=|C_{2}|=\cdots=|C_{c}|$ according to the Cauchy-Schwarz Inequality. A higher imbalance factor indicates a more uneven distribution of clusters. When most objects are concentrated in a few large clusters, the imbalance factor increases significantly, which is undesirable for our task. Overall, our goal is to achieve high-quality clustering results characterized by higher $P(C)$ and lower $\text{IF}(C)$. In our index, with proper hyperparameters, our training process of the cluster classifier is able to obtain high-quality clusters. In the experiments, We study the quality of the clusters produced by our proposed index using these evaluation metrics (shown in Section V-E).
IV-D Procedures and Analyses of LIST

### 18 Mar 2024 v2

@@ 3.1 Preliminary
We consider a geo-textual object dataset $D$, where each geo-textual object $o\in D$ has a location description $o.loc$ and a textual description $o.doc$. The location description $o.loc$ is a two-dimensional GPS coordinate composed of latitude and longitude. The textual description $o.doc$ is a document that describes the object. The Top-k KNN Spatial-Keyword Query (TkQcongEfficientRetrievalTopk2009 is defined as follows.
@@
3.2 Data Analysis and Motivations
@@
Word Mismatch. As discussed in Section 1, TkQ congEfficientRetrievalTopk2009 ; DBLP:conf/icde/FelipeHR08 ; DBLP:conf/ssd/RochaGJN11 typically uses traditional information retrieval methods such as BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 to compute textual relevance. These models rely on exact word matching to compute textual relevance and thus suffer from the word mismatch issue, which reduces their effectiveness. We further illustrate this issue in the real-world dataset that contains query-object ground-truth pairs. The details on the data are given in Section 1. As shown in Figure 1(a), in the Beijing dataset, nearly 20% of the ground-truth query-object positive (relevant) pairs have no overlap of words. This indicates that objects relevant to the query may receive a low or even zero textual relevance from these models. For instance, given a search ‘nearby drugstore’, if the search engine employs the BM25 to determine textual relevance, a drugstore labeled as ‘pharmacy’ would get a textual relevance score of zero, although the two terms convey the same meaning.
@@
Spatial Relevance. TkQ congEfficientRetrievalTopk2009 ; liu2023effectiveness typically computes spatial relevance by $1-SDist(q.loc,o.loc)$. The implicit assumption behind this linear function is that the user’s geographical preference for geo-textual objects is a linear function of distance. However, this assumption was not examined previously and we found it does not hold on our real-life datasets. In Figure 1(b), we illustrate this issue by comparing the cumulative distribution function (CDF) of ground-truth positive (relevant) query-object pairs with the ‘Linear Distribution’, a linearly ascending hypothetical scenario that positive query-object pairs are uniformly distributed across the range $[0,1]$. This figure shows a sharp increase in the CDF of ground-truth positive pairs for $SDist(q.loc,o.loc)$ below 0.1, a pattern that greatly differs from the hypothetical scenario. This motivates us to design a new spatial relevance module to better fit the real-world scenario. An intuitive solution would be to employ an exponential function. However, our experiments indicate that this method is even less effective compared to the linear function (as to be shown in Section 5.5).
@@
Efficiency. To address the word mismatch issue, several recent studies liu2023effectiveness ; DBLP:conf/aaai/ZhaoPWCYZMCYQ19 ; DBLP:conf/sigir/YuanLLZYZX20 have utilized deep learning techniques to evaluate textual relevance, thereby enhancing ranking effectiveness. However, these methods rely on word embeddings and complex neural networks to compute textual relevance, resulting in high querying latency. For instance, DrW liu2023effectiveness identifies the top-k relevant words from each object’s description for each query word based on their word embeddings, and then aggregates the scores of these word pairs using an attention mechanism to compute the textual relevance. On the Geo-Glue dataset, which comprises 2.8 million objects, DrW takes more than 7 seconds to answer a query on average in our experiment, which aligns with the results reported in liu2023effectiveness . This makes them unsuitable as a retriever for practical geo-textual object retrieval applications, although they can be used as re-rankers for a small number of objects returned by a retriever.
Motivations of LIST. As discussed in the last paragraph, incorporating deep textual relevance into spatial keyword queries has presented significant efficiency challenges. However, as discussed in Section 2, the first category of deep text relevance model, namely word interaction based deep relevance models, such as DrW liu2023effectiveness , which rely on word interaction modules and word embeddings to compute textual relevance, are not efficient for online computation and retrieval by a specifically designed index. Therefore, developing a new lightweight yet effective relevance model and a new index to accelerate the top-k search process for spatial keyword queries based on the proposed deep relevance model is in demand. For the first objective, we develop a lightweight yet effective deep relevance model for embedding based spatial keyword queries.
@@
For the second objective, existing geo-textual indexes DBLP:journals/vldb/ChenCCJ21 developed for spatial keyword queries are designed based on traditional retrieval models such as BM25 and cannot be used to handle embedding based spatial keyword queries. Existing ANNS indexes, which are discussed in Section 2, do not consider spatial factors. Directly employing existing ANNS indexes for spatial keyword queries will result in severe degradation of effectiveness (shown in Section 5.2). Simply adapting existing ANNS indexes, such as the IVF index, to incorporate spatial factors will cause even more severe effectiveness degradation, as it is hard to manually set a weight to balance spatial and textual similarities (as to be shown in Section 5.2). It is still an open problem to design a new ANNS index to support embedding based spatial keyword queries.
3.3 Problem Statement
In this study, we aim to (1) design a lightweight yet effective relevance model, and (2) develop an index to expedite the top-k search process of the proposed relevance model.
4 Proposed Retriever (LIST)
@@
4.1 Overview
LIST is featured with a new relevance model and a novel machine learning based ANNS index. Notably, our proposed index is also applicable to any other textual relevance model that transforms query keywords and object text descriptions into two separate embeddings.
The workflow of LIST is illustrated in Figure 2, which has three phases: the training, indexing, and query phase. During the training phase, we first train our relevance model, and then train our index. During the indexing phase, each object is partitioned into one cluster. During the query phase, the trained index routes each query to either a single cluster or a subset of clusters that have the highest probabilities. Within these clusters, LIST returns $k$ objects with the highest scores ranked by the trained relevance model as the query result.
4.2 The Proposed Relevance Model
Textual Relevance Learning. Inspired by the dual-encoder model’s success in document retrieval DBLP:conf/emnlp/KarpukhinOMLWEC20 , we employ a dual-encoder module to encode query keywords and object description into two seperate embeddings, and calculate textual relevance by the inner product between the two embeddings. Compared with previous relevance models liu2023effectiveness that rely on word embeddings and complex word interaction functions discussed in Section 2, the dual encoder module is both efficient and effective.
The dual-encoder module comprises an object encoder $E_{o}$ and a query encoder $E_{q}$, which are pre-trained language models like Bert devlin2018bert . The encoder takes the textual content of the query or object as input, captures interactions between words by a transformer-based model, and finally utilizes the representation of the [CLS] token as the global semantic representation, which is a $d$-dimensional embedding. This process is formulated as:
@@ 4.2 The Proposed Relevance Model
where $E_{o}(\textperiodcentered;\theta_{o})$ denotes the object encoder parameterized with $\theta_{o}$ and $E_{q}(\textperiodcentered;\theta_{q})$ denotes the query encoder parameterized with $\theta_{q}$. Then the text relevance score is calculated by the inner product between the $q.emb$ and $o.emb$,
@@
Spatial Relevance Learning. As discussed in Section 3.2, it is essential to develop a more effective spatial relevance module. Therefore, we propose a new learning-based spatial relevance module to learn to estimate spatial relevance.
As discovered in Section 3.2, the user’s geographic preferences for geo-textual objects do not follow a linear pattern. Here, we consider that the user’s geographic preference for geo-textual objects increases as the distance decreases, i.e., users tend to visit nearby objects, and exhibits a stepwise decline as the distance increases. To explain this stepwise pattern, let us consider a scenario in which a customer wishes to purchase coffee from Starbucks. If the nearest Starbucks is very close, s/he would go and buy coffee. If the nearest Starbucks is a little far, s/he may hesitate and her/his intention of visiting the Starbucks would decrease. If the nearest Starbucks is very far away, s/he would give up this idea. Therefore, the new spatial relevance module is designed to be a learnable monotonically increasing step function, to align with this real-world characteristic.
Our proposed spatial relevance module takes $S_{in}=1-SDist(q.loc,o.loc)$ as input, consisting of a threshold array $\mathcal{T}\in\mathbb{R}^{t\times 1}$ and a learnable weight array $w_{s}\in\mathbb{R}^{1\times t}$. Here, $\mathcal{T}$ stores the threshold values that determine the transition points of the step function, i.e., the value exceeded by $S_{in}$ will trigger an increase of spatial relevance, which is used to ensure that the learned function is a step function. Specifically, $\mathcal{T}$ is structured as $\mathcal{T}[i]=\frac{i}{t}$, where $i\in[0,t]$ and $t$ is a hyperparameter to control the increment of the threshold value. For example, when $t=100$, $\mathcal{T}$ is $[0.0,0.01,\cdots,0.99,1.0]$. The learnable weight array $w_{s}$ determines the extent of the increase when the input $S_{in}$ reaches these threshold values, which are learned from the training data and used to estimate the spatial relevance. When the input $S_{in}$ exceeds the value of $\mathcal{T}[i]$, then the spatial relevance increases by $act(w_{s}[i])$, where $act$ is an activation function to ensure $act(w_{s}[i])$ remain non-negative. This process ensures that the output $SRel$ exhibits a step increase as the input $S_{in}$ increases. The learned spatial relevance is computed as below:
@@
Adaptive Weight Learning. The recent study liu2023effectiveness has shown the importance of weight learning in improving ranking effectiveness. To enhance ranking effectiveness, we propose an adaptive weight learning module, which aims to assign adaptive weights to textual and spatial relevance based on the query keywords. For instance, in the case of detailed query keywords, such as ‘White House, 1600 Pennsylvania Avenue NW, Washington’, giving greater weight to textual relevance better fits the real-world scenario. Hence, we employ a simple yet effective manner that directly utilizes an MLP layer to determine the weights based on the embedding of query keywords $q.emb$, which is formulated as follows:
@@
Training Strategy. To train our relevance model, we employ the contrastive learning strategy. Given a query $q_{i}$, the positive (relevant) geo-textual object $o_{i}^{+}$ is obtained by real-world ground-truth data (detailed in Section 5.1). As for the negative (irrelevant) geo-textual objects, following previous study liu2023effectiveness , in each training epoch, we randomly pick $b$ hard negative samples from the dataset. We optimize the loss function as the negative log-likelihood of the positive object:
@@
In addition, we also utilize the in-batch negatives strategy lin2021inbatch to further enhance training efficiency.
Figure 3: The illustration of the index.
4.3 The Proposed Index
In this section, our objective is to develop a new ANNS index that, for each embedding based spatial keyword query, can return a subset of objects that receive high scores from the proposed relevance model for the input query, thereby accelerating the top-k search process.
To build such an index, our proposed index adopts a machine learning based method. Without manually setting the weight between spatial and textual relevance, it learns to cluster relevant queries and objects together while separating the irrelevant ones. Specifically, as illustrated in Figure 3, it takes the geo-location and textual embeddings of objects or queries as input and employs a Multi-Layer Perceptron (MLP) to partition objects and route queries into $c$ clusters, and then searches for the top-k relevant objects within the routed cluster. To make the index both effective and efficient, we develop a novel pseudo-negative query-object pair generation method. These pseudo labels combined with the positive (relevant) query-object pairs ensure that relevant queries and objects are partitioned and routed into the same cluster while separating the irrelevant ones, thereby reducing the search space while maintaining effectiveness.
@@ 4.3 The Proposed Index
Here $x$ is the input feature vector for the cluster classifier, i.e., the MLP.
Cluster Classifier. Our cluster classifier is a lightweight Multi-Perception Layer (MLP). Given $x$ as the input, it produces the $c$-cluster probability distribution, defined as follows:
@@
where $c$ is a hyperparameter that indicates the desired number of clusters we aim to obtain, and $Prob$ represents the predicted $c$-cluster probability distribution. The $c$-cluster probability distribution of object $o$ is denoted as $Prob_{o}$ and that of query $q$ is represented as $Prob_{q}$.
Note that the cluster classifier, i.e., the MLP, is shared between queries and objects. This setting allows our index to learn the distribution of both queries and objects, subsequently grouping identified relevant queries and objects into the same cluster (as detailed later).
Pseudo-Label Generation. To train the cluster classifier, both positive and negative pairwise labels are required. We utilize the ground-truth query-object relevant label (e.g., click-through or human annotation, detailed in Section 5.1) as the positive pairwise label for training. However, we lack high-quality negative pairwise labels. Randomly selecting negative objects will lead to overfitting and all objects are grouped together (as detailed later). Therefore, we propose a novel pseudo pairwise negative label generation method.
@@
This adaptive pseudo-negative generation method draws inspiration from hard negative sample strategy DBLP:conf/naacl/QuDLLRZDWW21 ; renRocketQAv2JointTraining2021 . What sets our approach apart from existing studies DBLP:conf/naacl/QuDLLRZDWW21 ; renRocketQAv2JointTraining2021 is that we employ an adaptive manner to select hard negative samples, thereby controlling the difficulty level of the generated negative samples. This adjustment strikes the trade-off between the effectiveness and efficiency.
@@
On the other hand, when our index is trained on a set of harder negative objects, it might also exclude some positive ones and reduce its effectiveness. Thus, the choice of $neg_{start}$ strikes a trade-off between effectiveness and efficiency. This alleviates the lack of negative label issue and the skewed cluster distribution issue of existing techniques Hsu16_KCL ; Hsu18_L2C .
Training Strategy. Based on the hard negative objects and positive objects provided above, we employ the MCL loss function Hsu19_MCL to train the MLP, as described below:
@@
As described earlier, the MLP is shared between queries and objects. Through this training process, for positive query-object pairs, their $Prob_{o}$ and $Prob_{q}$ will have a similar distribution. Thereby, relevant pairs of query and object are more likely to be grouped into the same cluster while the irrelevant pairs are grouped into distinct clusters. For example, as illustrated in Figure 3, the object $o_{4}$ and query $q_{2}$ form a positive query-object pair, and they are expected to be grouped into the same cluster $C_{2}$.
Learning to Partition and Route. During the indexing phase, each object $o$ is partitioned to the cluster with the highest probability according to $Prob_{o}$. Once the partitioning is completed, the objects assigned to a cluster are stored in a corresponding list, which acts as an inverted file for these objects. Each object in this list is represented by a $d$-dimensional vector $o.emb$, and a geo-location $o.loc$, which will be utilized to calculate the relevance score for incoming queries. During the query phase, a given query $q$ is directed to the cluster that has the highest probability. Subsequently, the relevance scores between $q$ and all objects within the cluster are calculated and the top-k objects are selected as result objects.
@@
Insertion and Deletion Policy. When a new object comes, we convert it into an embedding using the trained relevance model and then assign it to specific clusters using the trained index. When an object is deleted, we simply remove it from the corresponding cluster. Note that most existing ANNS indexes are static luoSurveyDeepHashing2023 ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 . When inserted data significantly differs in distribution from the existing data or when insertions occur frequently, periodically rebuilding the index is necessary to maintain high accuracy. Our index follows this approach. Apparently, the time costs of insertion and deletion operations are negligible.
Cluster Evaluation. After the training phase, we employ the trained index to produce clusters $C=\{C_{1},C_{2},\\ \cdots,C_{c}\}$, where $C_{i}$ represents cluster $i$. We use validation queries to evaluate the quality of the clusters. The validation queries are fed into the trained index and routed to a cluster. $C_{i}^{q}$ denotes a list containing the validation queries routed to $C_{i}$, $C_{i}^{o}$ denotes a list containing objects partitioned to $C_{i}$, and ||||| | indicates the size of a list. We proceed to introduce two metrics to evaluate the quality of the clusters.
@@
In addition, we also take into account efficiency concerns. For this purpose, we introduce another metric, the Imbalance Factor (IF) johnson2019billion , which measures the degree of balance across all clusters, denoted as $\text{IF}(C)$. The $IF(C)$ is formulated as:
@@
where $\text{IF}(C)$ is minimized when $|C_{1}|=|C_{2}|=\cdots=|C_{c}|$ according to the Cauchy-Schwarz Inequality. A higher imbalance factor indicates a more uneven distribution of clusters. When most objects are concentrated in a few large clusters, the imbalance factor increases significantly, which is undesirable for our task. Overall, our goal is to achieve high-quality clustering results characterized by higher $P(C)$ and lower $\text{IF}(C)$. In our index, with proper hyperparameters, our training process of the cluster classifier is able to obtain high-quality clusters. In the experiments, We study the quality of the clusters produced by our proposed index using these evaluation metrics (shown in Section 5.5).
4.4 Procedures and Analyses of LIST